OpenStudy (anonymous):

One angle in a triangle has a measure that is three times as large as the smallest angle. The measure of the third angle is 20 degrees more than that of the smallest angle. Find the measure of the LARGEST angle. Answer: The LARGEST angle has a measure of

OpenStudy (anonymous):

Well the angles in a triangle add up to 180 degrees total, for starters.

OpenStudy (anonymous):

Yeah so what would the formula to solve it look like? would it be x+x+20=180 ?

OpenStudy (yoloshroom):

Not quite

OpenStudy (yoloshroom):

alright so, the equation is 3x+20+x=180 , do you know why?

OpenStudy (anonymous):

Because the one angle is 3 times the other?

OpenStudy (yoloshroom):

yes, and one also 20 more than the smallest.

OpenStudy (anonymous):

so x=4

OpenStudy (anonymous):

I mean 40

OpenStudy (yoloshroom):

Wait, i may have messed up , let me check again

OpenStudy (yoloshroom):

Actually i think i may be right XD, sorry it's a bit late

OpenStudy (anonymous):

so is the largest angle 60

OpenStudy (inkyvoyd):

no

OpenStudy (yoloshroom):

no no, the largest would be 3x wouldnt it? so it'd be 3(40) so 120, whereas the whole equation would be 120+40+20 = 180 right?

OpenStudy (inkyvoyd):

it can't be 60. think about it: with an equilateral triangle, each angle is 60 degrees. any other triangle will have angles larger than 60.

OpenStudy (anonymous):

Okay I think I get it now

Directrix (directrix):

Does anyone think this is correct? Work courtesy of @Ms_Worth Let the smallest angle = x Then the third angle is ( x + 20 ) The second angle is 3x Together they add up to 180 x + 3x + ( x + 20 ) = 180 5x + 20 = 180 5x = 160 x = 32 <-- x was already defined as the smallest angle The other angles are ( x + 20 ) = 32 + 20 = 52 3x = 3 ( 32 ) = 96 <-- largest angle Answer: the largest angle is 96 degrees CHECK 32 + 52 + 96 = 180